# A Landmark-Aware Visual Navigation Dataset

### v1

@@ 6. Conclusion
In this work, we introduce a the first Landmark-Aware Visual Navigation (LAVN) dataset with human point-click annotations, providing exploration waypoint and environment landmark labels that allow for supervised training for exploration and map building tasks. We provide trajectories in 300 virtual and real-world environments with RGB observations, depth images, odometry, and human point-clicks for navigation waypoint supervision and environment landmark supervision. With this dataset, it will be possible to create visual navigation models that exhibit human-centric exploration policies to improve model performance on the image-goal task, paving the way for autonomous embodied agents in real-world applications.

### v2

@@ V Conclusion
In this work, we introduce the first Landmark-Aware Visual Navigation (LAVN) dataset with human point-click annotations, providing exploration waypoint and environment landmark labels that allow for supervised training for exploration and map building tasks. We provide trajectories in 300 virtual and real-world environments with RGB observations, depth images, odometry, and human point-clicks for navigation waypoint supervision and environment landmark supervision. With this dataset, it will be possible to improve visual navigation models that exhibit human-centric exploration policies on the image-goal task, paving the way for autonomous embodied agents in real-world applications. Future work includes incorporating dynamic objects in the scene.
Acknowledgment
This research was supported by the National Science Foundation (NSF) NRT NRT-FW-HTF: Socially Cognizant Robotics for a Technology Enhanced Socienty (SOCRATES), No. DGE-2021628 and Grant Nos. CNS-2055520, CNS-1901355, CNS-1901133.